A manufacturing company leases a building for $100,000 per year for its manufacturing facilities. In addition, the machinery in this building is being paid in installments of $18,000 per year. Each unit of the product produced costs $14 in labor and $11 in material. The product can be sold for $43. Assume that the demand in 10,000 unit per year.
What is the annual profit? Choose the closest answer below.

A. 37,200
B. 86,800
C. 99,200
D. 74,400
E. 62,000
Answer #1

Total units produced = 10,000 units

Price per unit = $43

Calculate the total revenue -

Total revenue = Price per unit * Total units produced

Total revenue = $43 * 10,000 = $430,000

The total revenue is $430,000

Annual lease of building = $100,000

Annual instalment for machinery = $18,000

Labor cost per unit = $14

Total labor cost = Labor cost per unit * Units produced = $14 * 10,000 = $140,000

Material cost per unit = $11

Total material cost = Material cost per unit * Units produced = $11 * 10,000 = $110,000

Calculate the total cost -

Total cost = Annual lease of building + Annual instalment for machinery + Total labor cost + Total material cost

Total cost = $100,000 + $18,000 + $140,000 + $110,000 = $368,000

The total cost is $368,000

Calculate profit -

Profit = Total revenue - Total cost

Profit = $430,000 - $368,000 = $62,000

Thus,

The annual profit is $62,000.

Hence, the correct answer is the option (E).